OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 33 feet and the child ride has a radius of 22 feet, what is the ratio between the volumes of each ride? Answer 27:8 9:6 9:4 3:2

OpenStudy (anonymous):

What do you think is the answer and why?

OpenStudy (anonymous):

First off do you know the formula for finding the volume of a cone?

OpenStudy (anonymous):

Vcone = πr2h

OpenStudy (anonymous):

area of a circle is pi r^2 so pi r^2 h will be the volume of a cylinder

OpenStudy (anonymous):

that's the right volume formula. Next you already know the radius is 22 and 33 (these are your values for r) Assuming then that the H is the same for both you have 1/3 (r1)^2 * pi = 1/3 (r2)^2 * pi

OpenStudy (anonymous):

where r1 is the first radius and r2 is the second.

OpenStudy (anonymous):

1/3 (r1)^2 * pi = 1/3 (r2)^2 * pi (r1)^2 * : (r2)^2

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Fine simply find the volume for each one individually

OpenStudy (anonymous):

once you have done that come back and post what you got. Also when finding the volume assume that there is no height. Dont even bother trying to plug it in.

OpenStudy (anonymous):

oh lol i got what u were sayin

OpenStudy (anonymous):

so use this formula 1/3 r^2 pi